tan x =1/2 and the terminal side of the angle is in the II quadrant ,then cos x =?


1
Expert's answer
2021-08-06T10:18:23-0400

"tanx =\frac{1}{2} \implies x=tan^{-1}0.5\\=26.57^0\\\text{Since the terminal side of the angle is in the 2nd quadrant, we have that $x=180^0-26.57^0$}\\\implies x=153.43^0
\\\text{Hence $cosx=-0.8944$}"